OpenStudy (anonymous):

Nellie is buying bird seed to put in her bird feeders. She uses about 10 pounds of bird seed each week, but has room to store any extra that she purchases.The bird seed comes in 3 different sizes as given below: 3.5 pounds for $2.52 7.5 pounds for $5.40 11 pounds for $7.92 Nellie is trying to be a smart shopper and get the best deal on the bird seed. Which size bag of bird seed should Nellie buy? a. Nellie should buy the 3.5 pound bags because they have the cheapest unit price b. Nellie should buy the 7.5 pound bags because they have the cheapest unit price c. Nellie should buy the 11 pound ba

OpenStudy (anonymous):

the answer c above isnt complete, could u complete it please?

OpenStudy (anonymous):

work out the price of 1 pound for each deal

OpenStudy (anonymous):

with the 3.5 pounds bag you do the following price of 1 pound = 2.52/3.5 = $0.72 with the 7.5 pounds bag price of 1 pound = 5.40/7.5 = $0.72 same as above one with the 11 pound bag price of 1 pound = ... can u do this one?

OpenStudy (anonymous):

i think there are more answers to follow, can u check please ur question above, (c) isn't ending completely and is there a (d) answer too?